OpenStudy (precal):

Differential equations - separable dy/dx=(1+lnx)y and if y=1 when x=1, then y=?

OpenStudy (precal):

when I separate them I get dy/y=(1+lnx)dx I don't know how to integrate the right side

OpenStudy (nincompoop):

As separate if want

OpenStudy (precal):

not sure I understand your statement

OpenStudy (nincompoop):

As. Separate equation

OpenStudy (nincompoop):

Lorry my tablet is not responding well

OpenStudy (precal):

that's ok, I am on my ipad. I am having issues as well

OpenStudy (precal):

what is the Q for? Boy is this website different

OpenStudy (superdavesuper):

\[\frac{ 1 }{y }dy = \left( 1 + lnx \right)dx = dx + lnx dx\] so integrating both sides \[\int\limits_{}^{}\frac{ dy}{ y } = \int\limits_{}^{}dx +\int\limits_{}^{}\ln(x)dx\] u can continue from here. good luck :)

OpenStudy (precal):

thanks

OpenStudy (precal):

This is an integration by parts problem. No wonder I forgot how to do it.